NOVEL RARE GAS CHEMISTRY

\underline{MARKKU RASANEN}, MIKA PETTERSSON, JAN LUNDELL, and LEONID KHRIACHTCHEV, Laboratory of Physical Chemistry, P.O.Box 55 (A.I.Virtasen aukio 1), FIN-00014 University of Helsinki, FINLAND.

UV-Photolysis of rare gas matrices doped with hydrogen halides yield H and Y (Y=halogen) atoms, and further excitation of the charge-transfer transition of XY pairs (X=Ar, Kr, Xe) results in mobilised holes. Permanent trapping of the holes may occur for instance in centers of type (XHX)^{+}. Annealing of the UV-irradiated H-containing matrices results in extremely strong IR-absorptions, which could by doping experiments be shown to belong to neutral species of type HXY. The novel rare-gas containing species found so far are HXeCl, HXeBr, HXeI, HKrCl and HXeH. All these species are strongly polar and their common structure is (HX)^{+}Y^{−}. In order to learn more about the potential surface, detailed spectroscopic studies are performed on HXeI. This species can be photodecomposed into \underline{neutral atoms} by infrared irradiation in the 3 μm region, and recovered quantitatively thermally at temperatures below 30 K. An Arrhenius analysis propose for the barrier from neutral atoms to HXeI a value of 700 cm^{−1}. We suggest that this barrier originates from an avoided crossing between two singlet surfaces limiting either to separated atoms H + Xe + I or to ionic species (HXe)^{+} + I^{−}.

References:

Neutral rare-gas containing charge-transfer molecules in solid matrices. I. HXeCl, HXeBr, HXeI and HKrCl in Kr and Xe. M.Pettersson, J.Lundell, M.Räsänen, J.Chem.Phys. 102 (1995) 6423-31.

Neutral rare gas containing charge-transfer molecules in solid matrices. II. HXeH, HXeD and DXeD in Xe. M.Pettersson, J.Lundell, M.Räsänen, J.Chem.Phys. 103 (1995) 205-10.

The mechanism of formation and IR-induced decomposition of HXeI in solid Xe. M.Pettersson, J.Nieminen, L.Khriachtchev, M.Räsänen, J.Chem.Phys. to be submitted.